### v1

@@ 1.2 HST Archival Detection
Hubble Space Telescope (HST) imaging of the cluster field reveals a clear early rest-frame far ultra-violet (FUV; Å) detection of SN Eos. This imaging was acquired on 2024 March 11, 13, 15, 17, 18 as part of an HST SNAP survey program to identify lensed stars [PID: 17504, kelly_snap_2023] in the WFC3/UVIS F814W and WFC3/IR F110W filters. The total exposure times for these two filters on each day were, in chronological order and for F814W & F110W respectively, [(2220 s, 1068 s), (1110 s, 609 s), (1110 s, 459 s), (880 s, 369 s), (880 s, 369 s)]222The fifth epoch of the WFC3/UVIS F814W data appears to have suffered a guide-star failure, and thus we do not include it in the analysis.. In a similar manner as for JWST, we process the HST data from the level 2 MAST products. For F814W, we use the products that include the pixel-based corrections for charge transfer efficiency (CTE). For both filters and all epochs (Table 1), we again drizzle exposures using the grizli pipeline [brammer_grizli_2023], following the approach used in the public DJA reduction. We use the native pixel scales for these products as well, corresponding to /pix for F814W and /pix for F110W.
@@ 1.3 VLT MUSE Archival Detection
Further archival Very Large Telescope (VLT) Multi Unit Spectroscopic Explorer (MUSE) data was retrieved from the ESO data archive that was taken on 2015 June 12 and July 17 – as a series of six exposures, each of 1462 s – thus preceding the JWST observation of images 101.1-2 (see Figure 1) of SN Eos by over a decade, or around 20 rest-frame months (data and analysis presented in [ciocan_muse_cluster_2021]). The locations of the SN appear coincident with the detection of an isolated emission line that corresponds to Ly emission and was perviously unreported. These detections are at a significance of 7.9, 12.0, 6.4, and 5.8 for images 101.1-4 respectively. Due to these data being taken well before the SN exploded in these images, this line is associated with the host galaxy of SN Eos and is also seen at two of the other host galaxy image locations (shown in Figure 3). A further dedicated analysis on the properties of this host will be presented in Y. Asada et al. (in prep.).
@@ 1.4 JWST DDT Observations
Spectrophotometric observations with JWST were executed on 2025 October 8 (DDT PID: 9493, PI: Coulter). This included a NIRSpec PRISM/CLEAR spectroscopic observation of both visible images 101.1 and 101.2 (with 10,504 s of exposure each), as well as follow-up NIRCam imaging (with 805 s of exposure) in the F277W, F356W, and F444W bands, each paired with the F070W SW filter (in an attempt to detect the underlying host). The NIRCam imaging was reduced using the standard VENUS procedure outlined in Section 1.1, while the NIRSpec spectra were reduced using the tools included in the msaexp package [brammer_msaexp_2023].
@@ 1.5 Photometry
We use STARRED [Michalewicz2023, Millon_2024] to measure HST photometry of SN Eos using the final drizzled products described in Section 1.1, which is a Python package for scene modeling, PSF photometry optimized for light curve extraction. In STARRED, the background is modeled on a grid of pixels with half the pixel size of the original image and regularized using starlets [starck_starlet_2015]. We use the template imaging (Figure 6) to constrain this static background, and then STARRED takes as input all time-series images along with a PSF model for each epoch to measure the SN light curve. The PSF is reconstructed with STARRED for each filter and epoch using 8 non-saturated stars in the field of view, and then each HST band is processed independently.
@@
STARRED does not perform well when there is no template imaging, as there is a degeneracy between the background reconstructon and the point source flux [pierel_cosmology_2025]. Therefore for JWST, where no template exists, we rely on robust aperture photometry with pixel radii, implemented on the final drizzled products, and background estimation using an annulus with inner and outer radii defined by the default reported aperture corrections (also applied to the raw measured fluxes) for each instrument. As all HST and JWST images were converted to the same units (nJy) by the grizli pipeline, all flux measurements were simply converted to AB magnitudes directly with the corresponding zero-point of . All measured photometry, including upper-limits, are reported in Table 1.
@@
Columns are: HST/JWST Program ID, Modified Julian date, JWST/HST instrument, WFC3/NIRCam filter, photometry from lens image 101.1 and 101.2 for SN Eos in AB mag. Upper limits are 5.
1.6 Lens Model
We construct a lens model for the galaxy cluster MACS 1931.8-2635 using a Python-implemented, updated version of the parametric method by Zitrin2015 where more details are provided). The lens model and its implementation will be detailed in separate companion article (Allingham et al., in prep.), and here we summarize its main properties.
The lens model consists of a dark-matter component and a galaxy component. With the cluster appearing mostly relaxed, we model its dark-matter component with one large, smooth dark-matter halo, corresponding to a pseudo-isothermal elliptical mass distribution [PIEMD; Keeton2001models]. Individual cluster galaxies are selected via a multi-band selected red-sequence, and are modeled each using a dual pseudo-isothermal elliptical mass distribution [dPIE, see Eliasdottir2007arXiv0710.5636E]. The cluster galaxies are scaled according to the Faber-Jackson relation [FaberJackson], but the BCG, and its ellipticity, are allowed to be modified independently.
Our lens model is based on 51 multiple images of 19 background sources that are used as constraints, 7 of which were previously known, e.g. from the CLASH survey [PostmanCLASH2012, Zitrin2015], and 12 are newly identified with NIRCam in the VENUS program (including the images of the SN and the host). The parameter values are optimized via a function minimizing the distance between observed images and their predicted location. The parameter space exploration was performed through a MCMC with stochastic moves guided by relative likelihood gradients.
The final best-fit model, minimized using 40 constraints and 21 free parameters, presents a error between the model predictions and observations of . Owing to stringent constraints requested on the position of SN Eos, the two existing images are reproduced with a precision lower than the pixel resolution (). Due to the low input positional uncertainties, the resultant time delay and magnification uncertainties are most likely underestimated. Thus, we adopt a nominal 20% additional systematic uncertainty on both throughout. The derived magnifications and time delays, located at the model-predicted image positions, are presented in Table 2.
@@ 1.6 Lens Model
1.7 Fe II pEW Measurement
To measure the pseudo-equivalent width pEW of Fe II 5018, we first smooth the spectrum using the inverse-variance Gaussian algorithm from branch06. From this smoothed spectrum, we identify the local maxima in flux on either side of the absorption feature to act as the endpoints of the linear pseudo-continuum. The value of the pEW is determined within this region and the uncertainty is estimated via a simple Monte Carlo simulation, drawing simulated fluxes from the flux uncertainties at each wavelength, and varying the smoothing scale for the continuum definition. We find a Fe II pEW = , a detection (see Figure 2 and Figure 4). We caution that given the low resolution of the detector, blending of adjacent SN features may bias this measurement. We also caution that the phase of SN Eos where we make this measurement is very near the end of the plateau, where the ionization state is changing due to a drop in the characteristic plateau temperature of K [Martinez2022_CSP1]. Nonetheless, this pEW is broadly consistent with an O3N2-derived metallicity of Z⊙ when considering that the pEW monotonically increases with phase, and is consistent with the expectation given the host.
Furthermore, we apply the same methods to measure the pEW of Fe II 5169. This feature is more strongly detected with Fe II pEW = . While there may be biases associated with deriving these values from low-resolution spectra, and we cannot rule out a potential contribution from the host galaxy [O III], we note that the Fe II absorption complex appears quite similar in morphology to the other low-metallicity local SNe-II (blue curves, Figure 4).
1.8 Spectroscopic Matching
We compare the two spectra of SN Eos to an extended library of low-redshift SN templates [2025RNAAS...9...78M] using SNID-SAGE [snid_sage_2025], a python implementation of the original Supernova Identification (SNID) [2007ApJ...666.1024B] code. Both spectra matched best to the +112 and +103 day spectra of SN 1992H, a Type IIP (plateau) SN that occurred in NGC 5377. Further shared matches include the low-metallicity Type II SN 2015bs, and the Type IIP SN 2013fs.
To compare the spectra of the two images of SN Eos directly to one another, we first scale the brighter 101.2 spectrum to the 101.1 spectrum using a Markov Chain Monte Carlo (MCMC) fit assuming a Gaussian posterior and uninformed prior over the spectral range of 2.2 – 5.5 m. This gives a scale factor of , consistent with the lens model-predicted value of . Since the continuum emission is much brighter than the contribution from individual lines, this process yields identical results when fitting to regions of the spectra that lack emission and absorption features. As shown in Figure 2, most of the commonly shared features between the spectra of the two images appear identical. We do note, however, that there appears to be a significant deviation in the flux of the H emission feature after scaling. Integrating the spectral flux density of the two spectra from 3.97 to 4.13 m, propagating the scale factor uncertainty into the significance, we find that the H flux is brighter for image 101.1 than image 101.2. If we instead scale the two spectra based on the observed photometry, we can derive a scale factor of , in agreement with the value derived directly from the spectra and in strong agreement with the lens model prediction. Using this scale factor instead, the significance in the H difference is on the order of . We believe that this difference in H strength is most likely resulting from contamination by the host galaxy. Due to the host being resolved, the total magnification factor affecting each image is different. This difference in magnification thus results in unique amounts of host contamination in the H line for each image, resulting in inconsistent H line strengths of the combined SN + host spectra. Based on the agreement in all the other line features, we believe that only H is significantly contaminated by the host, and we find no other evidence that could point to an alternate explanation, such as a large time delay or reduction artifact. An in-depth discussion of the underlying host contamination will be presented in Y. Asada et al. (in prep.).
1.9 SN Eos Modeling
To estimate the explosion properties of SN Eos, we compare synthetic light curves with the photometry of SN Eos presented in Table 1, demagnified by a factor of . We followed the same procedures as our previous studies [coulter_discovery_2025, moriya_jades_2025] on high-redshift Type II SNe to obtain synthetic light curves. We took the red supergiant (RSG) SN progenitor models in [moriya_jades_2025] having 0.1 Z⊙, in agreement with both the metallicity inferred from the analysis of Galaxy Eos and directly from SN Eos itself. The RSG progenitors were computed by the stellar evolution code Modules for Experiments in Stellar Astrophysics (MESA) [paxton2011mesa] and synthetic light curves were obtained by the one-dimensional multi-frequency radiation hydrodynamics code STELLA [blinnikov2000]. In general, the shock breakout [2023PASJ...75..634M] luminosity in the entire early UV light curve is not expected to be as luminous as the plateau optical luminosity, as is observed in SN Eos. To power this early UV emission and match the model to observations, we embedded the explosion within a spherically symmetric, dense, and confined circumstellar medium [forster2018] that is extended to with an enhanced mass-loss rate of , and a wind velocity of .
This simplified CSM prescription does not account for the first two UV light curve points in F110W and F814W, which could be due to the shock cooling emission of a larger radii star, or interaction with an aspherical CSM [Singh2024_aspherical_CSM, Yang2025_aspherical_CSM]. However, due to the uncertainty in absolute magnification and lack of complimentary early-time optical data, we instead focus instead on connecting the rise in the later UV points to the optical data at days. Our approach reasonably captures the UV light curve slopes and relative luminosities, and results in a confined CSM mass of 0.2 M⊙. This mass is considered a lower limit, as the limited UV data prevent us from knowing whether this UV flux was near maximum during the HST window of observations. We obtained synthetic light curves in the observer frame by shifting synthetic spectral energy distributions to and convolving them with the filter response functions. Figure 5 presents an example of a resultant light-curve model for SN Eos. This iteration is an explosion of a 16 M⊙ RSG progenitor having a hydrogen-rich envelope mass of 10.7 M⊙ with a radius of 593 R⊙. It has the explosion energy of and the 56Ni mass of 0.1 M⊙. The 56Ni mass is relatively high when compared to a sample of local SNe II [2019A&A...628A...7A], but it is close to that estimated for SN 1987A [2014ApJ...792...10S]. Mixing of 56Ni into the hydrogen-rich envelope affects the light-curve evolution [2019MNRAS.483.1211K], and we calculate two complimentary models with no mixing and full mixing, to create an envelope that matched SN Eos’s observed color and late-time light curve behavior.
1.10 Low- Analog Comparisons
SN Eos is the farthest spectroscopically classified supernova ever discovered, and is consistent with having a very low metallicity. Thus, we compare it directly to local Universe, low-metallicity (“low-”) SNe II which have been suggested to be good analogs to SNe at high-, as well as to a best-matching solar metallicity SN as a control. For the low- sample, there have only been three well-studied SNe II at : SN 2015bs [anderson_2015bs_2018], SN 2017ivv [Gutierrez2020_2017ivv], and SN 2023ufx [tucker_ufx_2024, Ravi2025_ufx]. For our solar-metallicity control, we compare against SN 1992H [1996AJ....111.1286C], which was our best spectral template match (see Section 1.8). The three low- SNe exhibit characteristics that are distinct from their solar-metallicity counterparts. Each SN was associated with a very low mass host galaxy, consistent with their low metallicity, having optical peaks brighter than the average [ AB mag; Richardson2014_AbsMag_dist], as well as faster declining plateau phases ( vs days) than the bulk of the population reported in the literature [see the following non-exhaustive list Anderson2014_SNIIP_LCs, Martinez2022_CSP1, Martinez2022_CSP_physical_params, Martinez2022_SNII_diversity]. In the case of SN Eos, we find that our model light curve has properties that span this comparison sample, with our inferred peak magnitude being similarly brighter than average ( AB mag), however, with a long duration plateau more closely matching SN 1992H (as well as other typical, local SNe II [Martinez2022_CSP1]).
Figure 2 compares SN Eos to these four SNe, selected at a similar phases (98 days) to our spectral observation of SN Eos. The exception is SN 2015bs, which we plot at a phase of days from explosion, because the spectrum at days did not have high enough signal-to-noise to compare Fe II. Remarkably, we find that SN Eos overall compares most favorably to SN 1992H, especially in terms of its H and Na I doublet (rest-frame Å) features. The disparity in H between SN Eos and SN 2017ivv may be due to a difference in phase, however, the boxy shape of SN 2023ufx’s H profile may indicate further interaction with an extended medium at days [Dessart2022_SNII_Interaction, although see the discussion in tucker_ufx_2024]. Such features are not apparent in the spectrum of SN Eos. In the right-hand panel of Figure 2, we focus on the Fe II absorption feature of each SN, and find extremely good agreement with all three low- SNe in the pEW of Fe II (including with SN 2015bs despite only comparing the day spectrum), whereas the absorption for SN 1992H is much deeper. More detailed comparisons between these SNe and SN Eos will require additional data: photometric observations of SN Eos’s decline off of the plateau, as well as nebular spectra, will constrain the amount of 56Ni synthesized, and therefore the core mass of the progenitor. Combining this with the envelope mass estimate from the plateau duration will enable a more confident inference on the zero-age main sequence mass of the progenitor to SN Eos, potentially holding implications for the IMF at .

### v2

@@ 6.2 HST Archival Detection
Hubble Space Telescope (HST) imaging of the cluster field reveals a clear early rest-frame far ultra-violet (FUV; Å) detection of SN Eos. This imaging was acquired on 2024 March 11, 13, 15, 17, 18 as part of an HST SNAP survey program to identify lensed stars [PID: 17504, (?)] in the WFC3/UVIS F814W and WFC3/IR F110W filters. The total exposure times for these two filters on each day were, in chronological order and for F814W & F110W respectively, [(2220 s, 1068 s), (1110 s, 609 s), (1110 s, 459 s), (880 s, 369 s), (880 s, 369 s)]222The fifth epoch of the WFC3/UVIS F814W data appears to have suffered a guide-star failure, and thus we do not include it in the analysis.. In a similar manner as for JWST, we process the HST data from the level 2 MAST products. For F814W, we use the products that include the pixel-based corrections for charge transfer efficiency (CTE). For both filters and all epochs (table S1), we again drizzle exposures using the grizli pipeline (?), following the approach used in the public DJA reduction. We use the native pixel scales for these products as well, corresponding to /pix for F814W and /pix for F110W.
@@ 6.3 VLT MUSE Archival Detection
Further archival Very Large Telescope (VLT) Multi Unit Spectroscopic Explorer (MUSE) data was retrieved from the ESO data archive that was taken on 2015 June 12 and July 17 – as a series of six exposures, each of 1462 s – thus preceding the JWST observation of images 101.1-2 (see Main Text Figure 1) of SN Eos by over a decade, or around 20 rest-frame months [data and analysis presented in (?)]. The locations of the SN appear coincident with the detection of an isolated emission line that corresponds to Ly emission and was previously unreported. These detections are at a significance of 7.9 , 12.0 , 6.4 , and 5.8 for images 101.1-4 respectively. Due to these data being taken well before the SN exploded in these images, this line is associated with the host galaxy of SN Eos and is also seen at two of the other host galaxy image locations (shown in Main Text Figure 3) . A further dedicated analysis on the properties of this host will be presented in Y. Asada et al. (in prep.).
@@ 6.4 JWST DDT Observations
Spectrophotometric observations with JWST were executed on 2025 October 8 (DDT PID: 9493, PI: Coulter). This included a NIRSpec PRISM/CLEAR spectroscopic observation of both visible images 101.1 and 101.2 (with 10,504 s of exposure each), as well as follow-up NIRCam imaging (with 805 s of exposure) in the F277W, F356W, and F444W bands, each paired with the F070W SW filter (in an attempt to detect the underlying host). The NIRCam imaging was reduced using the standard VENUS procedure outlined in Supplementary Text section 6.1, while the NIRSpec spectra were reduced using the tools included in the msaexp package (?).
@@ 6.5 Photometry
We use STARRED (?, ?) to measure HST photometry of SN Eos using the final drizzled products described in Supplementary Text section 6.1, which is a Python package for scene modeling PSF photometry, optimized for light curve extraction. In STARRED, the background is modeled on a grid of pixels with half the pixel size of the original image and regularized using starlets (?). We use the template imaging (figure S1) to constrain this static background, and then STARRED takes as input all time-series images along with a PSF model for each epoch to measure the SN light curve. The PSF is reconstructed with STARRED for each filter and epoch using 8 non-saturated stars in the field of view, and then each HST band is processed independently.
@@
STARRED does not perform well when there is no template imaging, as there is a degeneracy between the background reconstructon and the point source flux (?). Therefore for JWST, where no template exists, we rely on robust aperture photometry with pixel radii, implemented on the final drizzled products, and background estimation using an annulus with inner and outer radii defined by the default reported aperture corrections (also applied to the raw measured fluxes) for each instrument. As all HST and JWST images were converted to the same units ( nJy) by the grizli pipeline, all flux measurements were simply converted to AB magnitudes directly with the corresponding zero-point of . All measured photometry, including upper-limits, are reported in table S1.
6.6 Lens Model
We construct a lens model for the galaxy cluster MACS 1931.8-2635 using a Python-implemented, updated version of the parametric method by [ (?), but see also (?, ?) where more details are provided]. The lens model and its implementation will be detailed in separate companion article (Allingham et al., in prep.), and here we summarize its main properties.
The lens model consists of a dark-matter component and a galaxy component. With the cluster appearing mostly relaxed, we model its dark-matter component with one large, smooth dark-matter halo, corresponding to a pseudo-isothermal elliptical mass distribution [PIEMD; (?)]. Individual cluster galaxies are selected via a multi-band selected red-sequence, and are modeled each using a dual pseudo-isothermal elliptical mass distribution [dPIE, see (?)]. The cluster galaxies are scaled according to the Faber-Jackson relation (?), but the brightest cluster galaxy (BCG), and its ellipticity, are allowed to be modified independently.
Our lens model is based on 51 multiple images of 19 background sources that are used as constraints, 7 of which were previously known, e.g. from the CLASH survey (?, ?), and 12 are newly identified with NIRCam in the VENUS program (including the images of the SN and the host). The parameter values are optimized via a function minimizing the distance between observed images and their predicted locations. The parameter space exploration was performed through an MCMC with stochastic moves guided by relative likelihood gradients.
The final best-fit model, minimized using 40 constraints and 21 free parameters, presents an error between the model predictions and observations of . Owing to stringent constraints requested on the position of SN Eos, the two existing images are reproduced with a precision lower than the pixel resolution (). Due to the low input positional uncertainties, the resultant time delay and magnification uncertainties are most likely underestimated. Thus, we adopt a nominal 20% additional systematic uncertainty on both throughout. The derived magnifications and time delays, located at the model-predicted image positions, are presented in table S2.
6.7 Fe II pEW Measurement
To measure the pseudo-equivalent width pEW of Fe II 5018, we first smooth the spectrum using the inverse-variance Gaussian algorithm from (?). From this smoothed spectrum, we identify the local maxima in flux on either side of the absorption feature to act as the endpoints of the linear pseudo-continuum. The value of the pEW is determined within this region and the uncertainty is estimated via a simple Monte Carlo simulation, drawing simulated fluxes from the flux uncertainties at each wavelength, and varying the smoothing scale for the continuum definition. We find a Fe II pEW = , a detection (see Main Text Figure 2 and Figure 4) . We caution that given the low resolution of the detector, blending of adjacent SN features may bias this measurement. We also caution that the phase of SN Eos where we make this measurement is very near the end of the plateau, where the ionization state is changing due to a drop in the characteristic plateau temperature of K (?). Nonetheless, this pEW is broadly consistent with an O3N2-derived metallicity of Z⊙ when considering that the pEW monotonically increases with phase, and is consistent with the expectation given the host luminosity.
Furthermore, we apply the same methods to measure the pEW of Fe II 5169. This feature is more strongly detected with Fe II pEW = . While there may be biases associated with deriving these values from low-resolution spectra, and we cannot rule out a potential contribution from the host galaxy [O III], we note that the Fe II absorption complex appears quite similar in morphology to the other low-metallicity local SNe-II (blue curves, Main Text Figure 4).
6.8 Spectroscopic Matching
We compare the two spectra of SN Eos to an extended library of low-redshift SN templates (?) using SNID-SAGE (?), a Python implementation of the original Supernova Identification [SNID, (?)] code. Both spectra matched best to the +112 and +103 day spectra of SN 1992H, a Type IIP (plateau) SN that occurred in NGC 5377. Further shared matches include the low-metallicity Type II SN 2015bs, and the Type IIP SN 2013fs.
To compare the spectra of the two images of SN Eos directly to one another, we first scale the brighter 101.2 spectrum to the 101.1 spectrum using a Markov Chain Monte Carlo (MCMC) fit assuming a Gaussian posterior and uninformed prior over the spectral range of 2.2 – 5.5 m. This gives a scale factor of , consistent with the lens model-predicted value of . Since the continuum emission is much brighter than the contribution from individual lines, this process yields consistent results when fitting to regions of the spectra that lack emission and absorption features. As shown in Main Text Figure 2 , most of the commonly shared features between the spectra of the two images appear identical. We do note, however, that there appears to be a significant deviation in the flux of the H emission feature after scaling. Integrating the spectral flux density of the two spectra from 3.97 to 4.13 m, propagating the scale factor uncertainty into the significance, we find that the H flux is brighter for image 101.1 than image 101.2. If we instead scale the two spectra based on the observed photometry, we can derive a scale factor of , in agreement with the value derived directly from the spectra and in strong agreement with the lens model prediction. Using this scale factor instead, the significance in the H difference is on the order of . We believe that this difference in H strength is most likely resulting from contamination by the host galaxy. Due to the host being resolved, the total magnification factor affecting each image is different. This difference in magnification thus results in unique amounts of host contamination in the H line for each image, resulting in inconsistent H line strengths of the combined SN + host spectra. Based on the agreement in all the other line features, we believe that only H is significantly contaminated by the host, and we find no other evidence that could point to an alternate explanation, such as a large time delay or reduction artifact. An in-depth discussion of the underlying host contamination will be presented in Y. Asada et al. (in prep.).
6.9 SN Eos Modeling
To estimate the explosion properties of SN Eos, we compare synthetic light curves with the photometry of SN Eos presented in table S1, demagnified by a factor of . We followed the same procedures as our previous studies (?, ?) on high-redshift Type II SNe to obtain synthetic light curves. We took the red supergiant (RSG) SN progenitor models in (?) having 0.1 Z⊙, in agreement with both the metallicity inferred from the analysis of the LAE host and directly from SN Eos itself. The RSG progenitors were computed by the stellar evolution code Modules for Experiments in Stellar Astrophysics (MESA) (?) and synthetic light curves were obtained by the one-dimensional multi-frequency radiation hydrodynamics code STELLA (?). In general, the shock breakout (?) luminosity in the entire early UV light curve is not expected to be as luminous as the plateau optical luminosity, as is observed in SN Eos. To power this early UV emission and match the model to observations, we embedded the explosion within a spherically symmetric, dense, and confined circumstellar medium (?) that is extended to with an enhanced mass-loss rate of , and a wind velocity of .
This simplified CSM prescription does not account for the first two UV light curve points in F110W and F814W, which could be due to the shock breakout emission of a larger radii star, or interaction with an aspherical CSM (?, ?). However, due to the uncertainty in absolute magnification and lack of complimentary early-time optical data, we instead focus instead on connecting the rise in the later UV points to the optical data at days. Our approach reasonably captures the UV light curve slopes and relative luminosities, and results in a confined CSM mass of 0.2 M⊙. This mass is considered a lower limit, as the limited UV data prevent us from knowing whether this UV flux was near maximum during the HST window of observations. We obtained synthetic light curves in the observer frame by shifting synthetic spectral energy distributions to and convolving them with the filter response functions. Main Text Figure 5 presents an example of a resultant light-curve model for SN Eos. This iteration is an explosion of a 16 M⊙ RSG progenitor having a hydrogen-rich envelope mass of 10.7 M⊙ with a radius of 593 R⊙. It has the explosion energy of and the 56Ni mass of 0.1 M⊙. The 56Ni mass is relatively high when compared to a sample of local SNe II (?), but it is close to that estimated for SN 1987A (?). Mixing of 56Ni into the hydrogen-rich envelope affects the light-curve evolution (?), and we calculate two complimentary models with no mixing and full mixing, to create an envelope that matched SN Eos’ observed color and late-time light curve behavior.
6.10 Low- Analog Comparisons
SN Eos is the farthest spectroscopically classified supernova ever discovered, and is consistent with having a very low metallicity. Thus, we compare it directly to local Universe, low-metallicity (“low-”) SNe II which have been suggested to be good analogs to SNe at high-, as well as to a best-matching solar metallicity SN as a control. For the low- sample, there have only been three well-studied SNe II at : SN 2015bs (?), SN 2017ivv (?), and SN 2023ufx (?, ?). For our solar-metallicity control, we compare against SN 1992H (?), which was our best spectral template match (see Supplementary Text section 6.8). The three low- SNe exhibit characteristics that are distinct from their solar-metallicity counterparts. Each SN was associated with a very low mass host galaxy, consistent with their low metallicity, having optical peaks brighter than the average [ AB mag; (?)], as well as faster declining plateau phases ( vs days) than the bulk of the population reported in the literature [see the following non-exhaustive list, (?, ?, ?, ?)]. In the case of SN Eos, we find that our model light curve has properties that span this comparison sample, with our inferred peak magnitude being similarly brighter than average ( AB mag), however, with a long duration plateau more closely matching SN 1992H [as well as other typical, local SNe II (?)].
Main Text Figure 2 compares SN Eos to these four SNe, selected at a similar phases ( 98 days) to our spectral observation of SN Eos. The exception is SN 2015bs, which we plot at a phase of days from explosion, because the spectrum at days did not have high enough signal-to-noise to compare Fe II. Remarkably, we find that SN Eos overall compares most favorably to SN 1992H, especially in terms of its H and Na I doublet (rest-frame Å) features. The disparity in H between SN Eos and SN 2017ivv may be due to a difference in phase, however, the boxy shape of SN 2023ufx’s H profile may indicate further interaction with an extended medium at days[ (?), although see the discussion in (?)]. Such features are not apparent in the spectrum of SN Eos. In the right-hand panel of Main Text Figure 2 , we focus on the Fe II absorption feature of each SN, and find extremely good agreement with all three low- SNe in the pEW of Fe II (including with SN 2015bs despite only comparing the day spectrum), whereas the absorption for SN 1992H is much deeper. More detailed comparisons between these SNe and SN Eos will require additional data: photometric observations of SN Eos’ decline off of the plateau, as well as nebular spectra, will constrain the amount of 56Ni synthesized, and therefore the core mass of the progenitor. Combining this with the envelope mass estimate from the plateau duration will enable a more confident inference on the zero-age main sequence mass of the progenitor to SN Eos, potentially holding implications for the IMF at .